# PhysicsSolver: Transformer-Enhanced Physics-Informed Neural Networks for Forward and Forecasting Problems in Partial Differential Equations

### v1

@@ Abstract
Time-dependent partial differential equations are a significant class of equations that describe the evolution of various physical phenomena over time. One of the open problems in scientific computing is predicting the behaviour of the solution outside the given temporal region. Most traditional numerical methods are applied to a given time-space region and can only accurately approximate the solution of the given region. To address this problem, many deep learning-based methods, basically data-driven and data-free approaches, have been developed to solve these problems. However, most data-driven methods require a large amount of data, which consumes significant computational resources and fails to utilize all the necessary information embedded underlying the partial differential equations (PDEs). Moreover, data-free approaches such as Physics-Informed Neural Networks (PINNs) may not be that ideal in practice, as traditional PINNs, which primarily rely on multilayer perceptrons (MLPs) and convolutional neural networks (CNNs), tend to overlook the crucial temporal dependencies inherent in real-world physical systems. We propose a method denoted as PhysicsSolver that merges the strengths of two approaches: data-free methods that can learn the intrinsic properties of physical systems without using data, and data-driven methods, which are effective at making predictions. Extensive numerical experiments have demonstrated the efficiency and robustness of our proposed method. We provide the code at https://github.com/PhysicsSolver.
@@ 3 Related Works
Traditional Numerical Methods As a fundamental scientific problem, obtaining analytical solutions for partial differential equations is often difficult. Consequently, PDEs are usually discretized into meshes and solved using numerical methods in practice, such as finite difference method[21], finite element methods [8], spectral methods [19], etc. However, these numerical methods usually take a few hours or even days for complex structures [22] and perform badly on the forecasting problem.
@@
Data-free deep learning Method One of the most popular data-free deep learning methods is PINNs. This approach formalizes the constraints of PDEs, including the equations themselves, as well as initial and boundary conditions, as the objective functions within deep learning models [17, 28]. During the training process, the outputs of these models gradually align with the PDE constraints, allowing them to effectively approximate the solutions to the PDEs. However, this method mainly relies on convolutional neural networks, ignoring the crucial temporal dependencies inherent in practical physics systems, making it challenging to forecast solutions outside the given training grids. Also, directly using PINNs will fail to learn the solution of PDEs in some complex scenarios. For instance, [10, 14] propose the Asymptotic-Preserving PINNs to solve the difficulties caused by multiscale equations.
@@
Operator Learning-Based Methods Operator learning-based methods constitute a significant class within data-driven deep learning approaches, which involve training neural operators to approximate the input-output relationships in tasks governed by partial differential equations. This method can be applied to many physical scenarios, such as predicting future fluid behaviour based on past observations or estimating the internal stress in solid materials [15]. Some of the most well-known models in this area include the Fourier Neural Operator [12] and its variants, such as those described in [13] and [16]. There have been some works [27] focusing on forecasting the dynamics in PDEs. However, they often require a large volume of data and mostly they often ignore utilizing the intrinsic physical mechanism inside the PDEs.
@@
Transformer-Based Models The Transformer model [23] has garnered considerable attention for its capability to capture long-term dependencies, resulting in substantial advancements in natural language processing tasks [11]. Additionally, Transformers have been adapted for use in various other fields, including computer vision, speech recognition, and time-series analysis [5, 7, 25]. However, there are few researches on the application of transformers in solving PDEs. Recent researchers have used transformer to learn the solution of given PDEs [4, 26, 29]. However, the combination of PINNs and transformer hasn’t been fused that well and forecasting tasks in PDEs have not been studied enough.
@@ 5 Numerical Examples
Loss Design. For most of our experiments, we consider the spatial and temporal domains to be $[0,2\pi]$ and $[0,1]$ respectively. We choose the collocation points $\{(t_{i},x_{i})\}$ for $u(t,x)$ in the following way. For spatial points $x_{i}$, we select the interior points evenly spaced in $[0,2\pi]$. For temporal points $t_{i}$, we select $20$ interior points evenly spaced in the range $[0,1]$. We use the tensor product grid for the collocation points. For simplicity, the weights parameters in equation (10) are set to $(w_{1},w_{2},w_{3})=(1,1,1)$. The penalty parameters in equation (12) are set to $(\lambda_{1},\lambda_{2})=(1,1)$. Testing Settings. In terms of evaluation for two main tasks: the forward problem, and the forecasting problem, we adopted commonly used metrics in related works [10, 29], including the relative $\ell^{2}$ or $\ell^{\infty}$ error between the solution approximations $u^{\text{NN}}(t,x,\beta)$ and reference solutions $u^{\text{ref}}(t,x_{j},\beta)$, with the relative $\ell^{2}$ error defined by:
@@ 5.1 Linear Convection Equation
Forward Problem In the forward problem, we investigate the performance of different methods for the inference of the solution. Figure 2 shows the comparison for forecasting the solution between different methods. We can find that PINNs fails in capturing the solution with $t$ increases, while PINNsFormer and PhysicsSolver can capture the global solution well. More details can be found in Table 1, which shows that PhysicsSolver performs relatively better than PINNsFormer in the forward problem.
Figure 2: Forward problem for the convection equation.
@@
Table 1: Relative $l^{2}$ errors.
Single-step Forecasting Problem In the single-step forecasting problem, we aim to forecast the solution at the next time step. Figure 3 compares the solutions obtained by different methods at the last and second-to-last time points. It can be concluded that the Extrapolation method fails to accurately predict the future solution when the solution has relatively significant changes over time. More details can be found in Table 2. We can observe that PhysicsSolver outperforms the other three methods, thanks to its ability to learn a better mapping through the data system, which enhances its predictive capability.
Figure 3: Forecasting problem for convection equation.
@@
Table 2: Relative $l^{2}$ errors.
@@ 5.2 Reaction Equation
Forward Problem In the forward problem, we investigate the performance of different methods for the inference of the solution. Figure 4 shows the comparison for forecasting the solution between different methods. We can find that PINNs fails to capture the solution with $t$ increases, while PINNsFormer and PhysicsSolver can capture the global solution well. More details can be found in Table 3, it shows that PhysicsSolver performs much better than PINNsFormer in the forward problem thanks to its ability to learn a better mapping through the data system, which can improve the accuracy in the forward problem.
Figure 4: Forward problem for the convection equation.
@@
Error $9.803\times 10^{-1}$ $4.563\times 10^{-2}$ $\mathbf{5.550\times 10^{-3}}$
Table 3: Relative $l^{2}$ errors.
Single-step Forecasting Problem In the single-step forecasting problem, we aim to forecast the solution at the next time step. Figure 5 compares the solutions obtained by different methods at the last and second-to-last time points. It can be concluded that the PINNs method fails to predict the future solution accurately. However, the Extrapolation method works well when the solution has relatively small changes over time. More details can be found in Table 4. We can observe that PhysicsSolver outperforms the other three methods.
Figure 5: Forecasting problem for convection equation.
@@
Table 4: Relative $l^{2}$ errors.
5.3 Heat Diffusion PDE
@@
Forward Problem In the forward problem, we investigate the performance of different methods for the inference of the solution. Figure 6 shows the comparison for forecasting the solution between different methods. We can find that both the three methods can capture the solution as $t$ increases, while PINNsFormer and PhysicsSolver can capture the global solution better. More details can be found in Table 5, which shows that PhysicsSolver performs much better than PINNsFormer in the forward problem thanks to its ability to learn a better mapping through the data system, which improves the accuracy in the forward problem.
Figure 6: Forward problem for the heat equation.
@@
Table 5: Relative $l^{2}$ errors.
Single-step Forecasting Problem Similarly, we aim to forecast the solution at the next time step. Figure 7 compares the solutions obtained by different methods at the last and second-to-last time points. It can be concluded that the PINNs fails to accurately predict the future solution because it can not learn the time-dependent relationship between different solutions. The Extrapolation method fails to accurately predict the future solution when the solution has relatively significant changes over time. Also, PINNsformer can not capture the solution well since the model ignores the available data information. More details can be found in Table 6. We can observe that PhysicsSolver outperforms the other three methods.
Figure 7: Forecasting problem for convection equation.
@@
Table 6: Relative $l^{2}$ errors.
5.4 2D Navier-Stokes PDE
@@
Forward Problem In the forward problem, we investigate the performance of different methods for the inference of the solution for pressure $p$. Figure 8 shows the comparison for forecasting the solution between different methods. We can find that both the PINNs method and PINNsFormer method can not capture the solution well at different time stamps, while PhysicsSolver can capture the global solution better. More details can be found in Table 7, which shows that PhysicsSolver performs much better than PINNs and PINNsFormer in the forward problem thanks to its ability to learn a better mapping through the data system.
@@
Figure 8: Forward problem for the convection equation.
@@
Single-step Forecasting Problem In the single-step forecasting problem, we also aim to forecast the solution at the next time step. Figure 9 compares the solutions obtained by different methods at the last and second-to-last time points. It can be concluded that the PINNs fails to predict the future solution accurately since it can not learn the time-dependent relationship between different solutions well when the length of time is relatively large. At the same time, PINNsFormer can not capture the solution well since it overlooks the available data information. However, the Extrapolation method can accurately predict the future solution since the solution has relatively small changes over time. More details can be found in Table 8. We can observe that PhysicsSolver outperforms the other two methods and has a similar accuracy to the Extrapolation method.
Figure 9: Forecasting problem for convection equation.
@@ 6 Conclusions and future work
In this work, we propose an innovative transformer-based model called PhysicsSolver, which effectively solves both the forward and forecasting problems in PDE systems. Unlike previous approaches, our model can simultaneously learn intrinsic physical information and accurately predict future states by the ensemble of the physics system and the data system. We conduct extensive numerical experiments and the results show the superiority of our proposed method. For future work, we will also consider the application of different variants of PhysicsSolver for solving more complicated problems such as multi-scale kinetic systems.

### v2

@@ Abstract
Time-dependent partial differential equations are a significant class of equations that describe the evolution of various physical phenomena over time. One of the open problems in scientific computing is predicting the behaviour of the solution outside the given temporal region. Most traditional numerical methods are applied to a given time-space region and can only accurately approximate the solution of the given region. To address this problem, many deep learning-based methods, basically data-driven and data-free approaches, have been developed to solve these problems. However, most data-driven methods require a large amount of data, which consumes significant computational resources and fails to utilize all the necessary information embedded underlying the partial differential equations (PDEs). Moreover, data-free approaches such as Physics-Informed Neural Networks (PINNs) may not be that ideal in practice, as traditional PINNs, which primarily rely on multilayer perceptrons (MLPs) and convolutional neural networks (CNNs), tend to overlook the crucial temporal dependencies inherent in real-world physical systems. We propose a method denoted by PhysicsSolver that merges the strengths of two approaches: data-free methods that can learn the intrinsic properties of physical systems without using data, and data-driven methods, which are effective at making predictions. Extensive numerical experiments have demonstrated the efficiency and robustness of our proposed method. We provide the code at https://github.com/PhysicsSolver.
@@ 3 Related Works
Traditional Numerical Methods. As a fundamental scientific problem, obtaining analytical solutions for partial differential equations is often difficult. Consequently, PDEs are usually discretized into meshes and solved using numerical methods in practice, such as finite difference methods [25], finite element methods [10], spectral methods [22], etc. However, these numerical methods usually take a few hours or even days for complex structures [28] and perform badly on the forecasting problem.
@@
Data-free deep learning Method. One of the most popular data-free deep learning methods is PINNs. This approach formalizes the constraints of PDEs, including the equations themselves, as well as initial and boundary conditions, as the objective functions within deep learning models [20, 34]. During the training process, the outputs of these models gradually align with the PDE constraints, allowing them to effectively approximate the solutions to the PDEs. However, this method mainly relies on convolutional neural networks, ignoring the crucial temporal dependencies inherent in practical physics systems, making it challenging to forecast solutions outside the given training grids. Also, directly using PINNs will fail to learn the solution of PDEs in some complex scenarios. For instance, [12, 16] propose the Asymptotic-Preserving PINNs to solve the difficulties caused by multiscale equations. Except the methods mentioned above, [7] proposed the PIKFNNs, which can improve the performance of the traditional PINNs by embedding PDE information into the activation function of neural networks.
@@
Operator Learning-Based Methods. Operator learning-based methods constitute a significant class within data-driven deep learning approaches, which involve training neural operators to approximate the input-output relationships in tasks governed by partial differential equations. This method can be applied to many physical scenarios, such as predicting future fluid behaviour based on past observations or estimating the internal stress in solid materials [17]. Some of the most well-known models in this area include the Fourier Neural Operator [14] and its variants, such as those described in [15] and [19]. There have been some works [33] focusing on forecasting the dynamics in PDEs. However, they often require a large volume of data and mostly they often ignore utilizing the intrinsic physical mechanism inside the PDEs.
@@
Transformer-Based Models. The Transformer model [29] has garnered considerable attention for its capability to capture long-term dependencies, resulting in substantial advancements in natural language processing tasks [13]. Additionally, Transformers have been adapted for use in various other fields, including computer vision, speech recognition, and time-series analysis [5, 9, 31]. So far, there are few researches on the application of transformers in solving PDEs. Recent researchers have used transformer to learn the solution of given PDEs [4, 32, 35]. However, the combination of PINNs and transformer hasn’t been fused that well and forecasting tasks in PDEs have not been studied enough.
Aside from the topics of solving PDEs, there are several other intriguing subjects worth exploring, such as the training challenges associated with PINNs. For example, [18] introduces a unified scalable framework designed to address the training difficulties encountered in PINNs and XPINNs, particularly for time-dependent PDEs. Additionally, [8] presents a pre-training PINN framework aimed at enhancing the convergence and accuracy of the standard PINNs method, while [23] offers notable theoretical advancements in applying deep learning to solve PDEs.
@@ 5 Numerical Examples
Loss Design. For most of our experiments, we consider the spatial and temporal domains to be $[0,2\pi]$ and $[0,1]$ respectively. We choose the collocation points $\{(t_{i},x_{i})\}$ for $u(t,x)$ in the following way. For spatial points $x_{i}$, we select the interior points evenly spaced in $[0,2\pi]$. For temporal points $t_{i}$, we select $20$ interior points evenly spaced in the range $[0,1]$. We use the tensor product grid for the collocation points. For simplicity, the weights parameters in equation (10) are set to $(w_{1},w_{2},w_{3})=(1,1,1)$. The penalty parameters in equation (12) are set to $(\lambda_{1},\lambda_{2})=(1,1)$.
Testing Settings. In terms of evaluation for two main tasks: the forward problem, and the forecasting problem, we adopted commonly used metrics in related works [12, 35], including the relative $\ell^{2}$ or $\ell^{\infty}$ error between the solution approximations $u^{\text{NN}}(t,x,\beta)$ and reference solutions $u^{\text{ref}}(t,x_{j},\beta)$, with the relative $\ell^{2}$ error defined by:
@@ 5.1 Linear Convection Equation
Forward Problem. In the forward problem, we investigate the performance of different methods for the inference of the solution. Figure 2 shows the comparison for forecasting the solution between different methods. We can find that PINNs fails in capturing the solution with $t$ increases, while PINNsFormer and PhysicsSolver can capture the global solution well. Furthermore, we show the error distribution across various methods for the convection equation in Figure 3. More details can be found in Table 1, which shows that PhysicsSolver performs relatively better than PINNsFormer in the forward problem.
Figure 2: Performances of methods to the forward problem for the convection equation.
Figure 3: Error distribution across various methods for the convection equation.
@@
Table 1: Relative $l^{2}$ errors for the forward problem.
Single-step Forecasting Problem. In the single-step forecasting problem, we aim to forecast the solution at the next time step. Figure 4 compares the solutions obtained by different methods at the last and second-to-last time points. It can be concluded that the Extrapolation method fails to accurately predict the future solution when the solution has relatively significant changes over time. More details can be found in Table 2. We can observe that PhysicsSolver outperforms the other three methods, thanks to its ability to learn a better mapping through the data system, which enhances its predictive capability.
Figure 4: The forecasting problem for the convection equation.
@@
Table 2: Relative $l^{2}$ errors for the forecasting problem.
@@ 5.2 Reaction Equation
Forward Problem. In the forward problem, we investigate the performance of different methods for the inference of the solution. Figure 5 shows the comparison for forecasting the solution between different methods. We can find that PINNs fails to capture the solution with $t$ increases, while PINNsFormer and PhysicsSolver can capture the global solution well. Furthermore, we show the error distribution across various methods for the reaction equation in Figure 6. More details can be found in Table 3, it shows that PhysicsSolver performs much better than PINNsFormer in the forward problem thanks to its ability to learn a better mapping through the data system, which can improve the accuracy in the forward problem.
Figure 5: Performances of methods to the forward problem for the reaction equation.
Figure 6: Error distribution across various methods for the reaction equation.
@@
Error $9.803\times 10^{-1}$ $4.563\times 10^{-2}$ $\mathbf{5.550\times 10^{-4}}$
Table 3: Relative $l^{2}$ errors for the forward problem.
Single-step Forecasting Problem. In the single-step forecasting problem, we aim to forecast the solution at the next time step. Figure 7 compares the solutions obtained by different methods at the last and second-to-last time points. It can be concluded that the PINNs method fails to predict the future solution accurately. However, the Extrapolation method works well when the solution has relatively small changes over time. More details can be found in Table 4. We can observe that PhysicsSolver outperforms the other three methods.
Figure 7: The forecasting problem for the reaction equation.
@@
Table 4: Relative $l^{2}$ errors for the forecasting problem.
5.3 Heat Diffusion Equation
@@
Forward Problem. In the forward problem, we investigate the performance of different methods for the inference of the solution. Figure 8 shows the comparison for forecasting the solution between different methods. We can find that both the three methods can capture the solution as $t$ increases, while PINNsFormer and PhysicsSolver can capture the global solution better. Furthermore, we show the error distribution across various methods for the heat equation in Figure 9. More details can be found in Table 5, which shows that PhysicsSolver performs much better than PINNsFormer in the forward problem thanks to its ability to learn a better mapping through the data system, which improves the accuracy in the forward problem.
Figure 8: Performances of methods to the forward problem for the heat equation.
Figure 9: Error distribution across various methods for the heat equation.
@@
Table 5: Relative $l^{2}$ errors for the forward problem.
Single-step Forecasting Problem. Similarly, we aim to forecast the solution at the next time step. Figure 10 compares the solutions obtained by different methods at the last and second-to-last time points. It can be concluded that the PINNs fails to accurately predict the future solution because it can not learn the time-dependent relationship between different solutions. The Extrapolation method fails to accurately predict the future solution when the solution has relatively significant changes over time. Also, PINNsformer can not capture the solution well since the model ignores the available data information. More details can be found in Table 6. We can observe that PhysicsSolver outperforms the other three methods.
Figure 10: Forecasting problem for the heat equation.
@@
Table 6: Relative $l^{2}$ errors for the forecasting problem.
5.4 2D Navier-Stokes Equations
@@
Forward Problem. In the forward problem, we investigate the performance of different methods for the inference of the solution for pressure $p$. Figure 11 shows the comparison for forecasting the solution between different methods. We can find that both the PINNs method and PINNsFormer method can not capture the solution well at different time stamps, while PhysicsSolver can capture the global solution better. Furthermore, we show the error distribution across various methods for the Navier-Stokes equation in Figure 12. More details can be found in Table 7, which shows that PhysicsSolver performs much better than PINNs and PINNsFormer in the forward problem thanks to its ability to learn a better mapping through the data system.
@@
Figure 11: Forward problem for the Navier-Stokes equation.
Figure 12: Error distribution across various methods for the Navier-Stokes equation.
@@
Single-step Forecasting Problem. In the single-step forecasting problem, we also aim to forecast the solution at the next time step. Figure 13 compares the solutions obtained by different methods at the last and second-to-last time points. It can be concluded that the PINNs fails to predict the future solution accurately since it can not learn the time-dependent relationship between different solutions well when the length of time is relatively large. At the same time, PINNsFormer can not capture the solution well since it overlooks the available data information. However, the Extrapolation method can accurately predict the future solution since the solution has relatively small changes over time. More details can be found in Table 8. We can observe that PhysicsSolver outperforms the other two methods and has a similar accuracy to the Extrapolation method.
Figure 13: Forecasting problem for the Navier-Stokes equation.
@@
5.5 Special Case: multi-step Forecasting Problem for the Reaction Equation
In the single-step forecasting problem, $\tilde{t}$ is a single time step, which is one time step later than previous time $t$ used for learning. In multi-steps forecasting problem, suppose the time steps involved to train are $\{t_{m}\}_{m=1}^{m_{1}}$, and the time steps to be predicted are denoted by $\{t_{m}\}_{m=m_{1}+1}^{m_{2}}$. There are several different methods for multi-steps forecasting, such as direct multi-step forecast strategy, recursive multi-step forecast strategy, hybrid forecast strategies, we refer the readers to [26, 27]. Here we study the direct multi-step forecast strategy for PINNs, PINNsFormer and PhysicsSolver, which means we directly generate the solutions on $\{t_{m}\}_{m=m_{1}+1}^{m_{2}}$ time steps by utilizing the pre-trained models. For the extrapolation method, one can apply the recursive extrapolation approach to obtain the solution on multi-time steps.
First, we investigate the performance of different methods for $\Delta t=0.1$. The comparison results can be found in Figure 14. Table 9 shows the detailed results which demonstrates that our proposed method outperforms other existing methods in terms of accuracy. We observe that our PhysicsSolver algorithm can achieve the relative $l^{2}$ error as small as $1.79\times 10^{-3}$.
Figure 14: Forecasting problem with $\Delta t=0.1$ across various methods for the reaction equation.
Method Extrapolation PINNs PINNsFormer PhysicsSolver
Error $5.015\times 10^{-3}$ $9.673\times 10^{-1}$ $6.450\times 10^{-2}$ $\mathbf{1.793\times 10^{-3}}$
Table 9: Relative $l^{2}$ errors for the forecasting problem with $\Delta t=0.1$.
Moreover, we investigate the performance of different methods for a larger time $\Delta t=0.5$. The comparison results are shown in Figure 15 and Table 10 for the relative $l^{2}$ errors of different approaches. We conclude that our PhysicsSolver significantly surpasses other methods in the long-term forecasting problem, reaching a level of accuracy as small as $\mathcal{O}(10^{-3})$.
Figure 15: Forecasting problem with $\Delta t=0.5$ across various methods for the reaction equation.
Method Extrapolation PINNs PINNsFormer PhysicsSolver
Error $7.639\times 10^{-2}$ $9.462\times 10^{-1}$ $7.785\times 10^{-2}$ $\mathbf{5.735\times 10^{-3}}$
Table 10: Relative $l^{2}$ errors for the forecasting problem with $\Delta t=0.5$.
@@ 6 Conclusions and future work
In this work, we propose an innovative transformer-based model called PhysicsSolver, which effectively solves both the forward and forecasting problems in PDE systems. Unlike previous approaches, our model can simultaneously learn intrinsic physical information and accurately predict future states by the ensemble of the physics system and the data system. We conduct extensive numerical experiments and the results show the superiority of our proposed method.
Limitations and future work: First, we lack of theoretical analysis such as the convergence of predicted solution to reference solution in our work, as in most literature. Also, due to data limitations in certain examples, predicting PDEs with large time instants across more cases poses a significant challenge. In the future work, we plan to extend our method and study multi-step forecasting problem in broader applications. As time evolves, the error of extrapolation procedure may increase, requiring our current method to incorporate more constraints in the design of neural networks. Moreover, we intend to explore different variants of PhysicsSolver in tackling more complicated problems such as multi-scale kinetic models.